Here is a balanced equation. Use it to answer the questions below.

2Pb(NO3)2
1
Expert's answer
2020-05-12T11:36:19-0400

2Pb(NO3)2 ---> 2PbO + 4NO2 + O2


4 moles of nitrogen dioxide will be produced from 2 moles of lead(ll) nitrate


0.5 mole of oxygen will be produced from 1 mole of lead(ll) nitrate


The mass of PbO obtained from 200g of Pb(NO3)2:

M (PbO) = 223.20

M (Pb(NO3)2) = 331.21

x = 200* 223.20 / 331.21 = 134.78 g


If 130g of PbO are actually obtained when the experiment is done in the laboratory, the percentage yield is:

130 / 134.78 = 0.965 or 96.5%